A 900 kg car has a net force of 2400 N acting on it. What is the acceleration of the car? Express your answer in meters per second squared and round to one decimal place
Answer #1

Newton's second law of motion is F = ma, or force is equal to mass times acceleration.

so F =ma

then acceleration a = F/m

given data

net force F= 2400 N

car has mass m= 900 kg

plugging the values in the formula

a = F/m = 2400N/900 kg = 2.67 m/s2

a =  2.67 m/s2

a = 2.7 m/s2

The acceleration of the car = a = 2.7 m/s2

if you find this helpful please up vote